OpenStudy (anonymous):

Essay: Show all work. A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x-8y ft. and the middle length side x+6y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out and express in simplest correct mathematical form, including units. Help on this, What is the algebraic expression?

OpenStudy (anonymous):

Area= 1/2 b h b= x-8y h= x+6y 1/2 (x-8y) (x+6y)

OpenStudy (anonymous):

1/2(x-8y) + 1/2(x+6y) 1/2x -4y + 1/2x+3y 1x-1y x-y

OpenStudy (anonymous):

A=1/2b*h not 1/2(b+h) so its: A=1/2(x-8y)(x+6y) A=1/2(x^2-2xy-48y^2) A=(1/2)x^2-xy-24y^2